Write the slope-intercept equation for the line with slope 4/5 and y-intercept (0, -1) and graph the line
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b where m is the slope and b the y-intercept.
slope-intercept equation for the line with slope 4/5 and y-intercept (0, -1)
y = (4/5)x - 1
Note: as y increases/decreases by 4, x increases/decreases by 5


Answer by Alan3354(69443)

It's y = mx + b
----------
y = (4/5)x - 1
To graph it, get the FREE software at
http://www.padowan.dk